# MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning

### v2

@@ Abstract
Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory cost. Previous LoRA-based approaches initialize the low-rank matrices with Gaussian distribution and zero values while keeping the original weight matrices frozen. However, the trainable model parameters optimized in an unguided subspace might interfere with the well-learned subspace of the pretrained weight matrices. In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principal singular components frozen. It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principal matrix contains important knowledge. The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principal matrix, thus the pretrained knowledge is expected to be well preserved. During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the labeled dataset. Extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks present the superior performance of our method.
MiLoRA: Harnessing Minor Singular Components for Parameter-Efficient LLM Finetuning
Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2,  Yun Chen1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology   3Tsinghua University
@@ 1 Introduction
Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024a) and question answering (Ivison et al., 2023). These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks. Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
@@
In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning. MiLoRA has a similar framework as LoRA but employs a different initialization schedule. Specifically, a weight matrix ${\bm{W}}$ is decomposed with the singular value decomposition (SVD) algorithm. Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principal matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values. We argue that the principal matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information. It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
Motivated by these observations, we freeze the principal matrix ${\bm{W}}_{p}$ and adapt the minor singular components during finetuning. The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$. Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principal matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge. To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros. In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals the frozen principal part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$. We conduct extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks. The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+1.4 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B). [^1] Our code and model will be made public.
@@ 2 Preliminaries
Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$. The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors. The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order.
Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as in Equation 1.
@@
where ${\bm{W}}^{(0)}$ and $\Delta{\bm{W}}$ refer to the pretrained weight and weight change, respectively. The $\alpha$ and $r$ are hyperparameters of scaling-factor and LoRA rank ($r\ll\min(m,n)$). During finetuning, the pretrained matrix ${\bm{W}}^{(0)}$ is kept frozen. It significantly diminishes the number of trainable parameters as both ${\bm{A}}$ and ${\bm{B}}$ matrices are low-rank. The ${\bm{B}}$ matrix is initialized with zero while ${\bm{A}}$ matrix adopts a random Gaussian distribution with zero mean value.
This initialization strategy ensures the $\Delta{\bm{W}}=0$ at the beginning of training. The LoRA method only modifies the linear matrices in the Transformer model. The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with full finetuning. However, the vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized. This might potentially detract from the pretrained knowledge encoded in the pretrained weight matrix.
@@ 3 Methodology
The decomposition can be also reformulated in a matrix form. The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principal and minor singular values, respectively. The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated similarly. The SVD of ${\bm{W}}$ can be expressed as:
@@
To make our method simple and training-efficient, we do not incorporate the regularization loss which imposes strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$. Our experiment results show that our scheme works quite well in balancing learning from the finetuning dataset and preserving pretrained knowledge.
@@ 4 Experiments
Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks. Results are cited from Liu et al. (2024b). Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
@@
Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. Results are cited from Yu et al. (2023).
@@ 4.1 Experiments on Large Language Model
We evaluate and compare MiLoRA with baselines on three different types of LLM downstream tasks.
1. Commonsense reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) and LLaMA3-7B (AI@Meta, 2024) on Commonsense170K (Hu et al., 2023). Eight commonsense reasoning datasets are used for evaluation, including BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). The task is formulated as a multiple-choice problem. We report accuracy (%) for all datasets on the best checkpoint chosen by the validation set loss.
2. Math reasoning: We finetune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains 395K samples augmented from the training set of GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
We use test sets of GSM8K and MATH for evaluation and report results on the last checkpoint. For each test set, we report the exact match ratio against the ground truth.
3. Instruction following: We follow the experiment setup in Wu et al. (2024a) and Wu et al. (2024b) to finetune LLaMA2-7B with Ultrafeedback (Cui et al., 2023). Following previous works, we use Alpaca-Eval v1.0 (Li et al., 2023a) for evaluation and report win-rate against text-davinci-003 using GPT4-0613 as the annotator (OpenAI, 2024).
@@
We use the same hyperparameter configurations as Hu et al. (2023) without tuning for all methods. Details can be found in Appendix A.2. We denote this hyperparameter setup as LLM-Adapters. By default, we use a rank of 32. As math reasoning has a large training dataset, we set the rank to 64.
Method Win-rate
GPT-4-0613 91.2
GPT-3.5-turbo-0613 91.8
LLaMA2-Chat-7B 87.3
LLaMA2-7B + LoRA 85.9
LLaMA2-7B + PiSSA 85.6
LLaMA2-7B + MiLoRA 87.3
Table 3: Instruction tuning evaluation results for instruction-tuned LLaMA2 7B with Alpaca-Eval v1.0. Results are cited from Ivison et al. (2023).
@@ Results
Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks. We directly use checkpoints from Liu et al. (2024a) to reproduce the results of baseline methods.
We report results in Table 1, Table 2 and 3 for commonsense reasoning, math reasoning and instruction following, respectively. For commonsense reasoning, we also include the ChatGPT baseline reported in Liu et al. (2024b) as a reference, which is obtained with gpt-3.5-turbo API using a zero-shot Chain of Thought (Wei et al., 2022). For math reasoning, we add the full finetuning results from Yu et al. (2023) as a reference. For instruction following, we include the results of GPT-4-0613, GPT-3.5-turbo-0613, and LLaMA2-Chat-7B from Ivison et al. (2023) as a reference.
As can be seen, MiLoRA consistently surpasses all baseline methods across different datasets and LLMs, indicating that MiLoRA is a highly effective finetuning method. Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B, and by 1.1 and 6.5 points on LLaMA3-8B. For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged EM points on LLaMA2-7B, respectively. However, it still underperforms full finetuning by an average of 2.5 EM score. This suggests that the PEFT methods still have room for improvement to fully match the performance of full finetuning. For instruction following, MiLoRA achieves the best results, surpassing LoRA and PiSSA by 1.4 and 2.7 in win-rate scores, respectively.
@@ 5.1 Comparison with More PEFTs
To further demonstrate the effectiveness of MiLoRA, we compare MiLoRA with additional popular baselines, such as rsLoRA Kalajdzievski (2023), LoRA+ Hayou et al. (2024), DoRA Liu et al. (2024b), AdaLoRA Zhang et al. (2023), LoRA-GA Wang et al. (2024b). We follow the math and code experiment settings in Wang et al. (2024b). Specifically, we finetune LLaMA2-7B with MiLoRA on MetaMathQA 100K and Code-Feedback 100K for one epoch using rank $8$ and compare MiLoRA with baseline results reported in Wang et al. (2024b). Table 5 shows the results. We report the training cost of all methods relative to LoRA, which requires 3.58 hours of training on the math task using a single NVIDIA-L40 GPU. As can be seen, MiLoRA consistently outperforms all baselines, improving the best PEFT baseline LoRA-GA by 1.1 and 4.6 points on GSM8K and Human-eval, respectively. The training efficiency of MiLoRA is also very high. The time for SVD decomposition in MiLoRA is less than six minutes and negligible when compared with the total training time. Therefore, the training time required of MiLoRA is consistent with that of LoRA, while DoRA and AdaLoRA require additional training time overhead.
@@
Table 5: Comparison with more baselines on math and code tasks. Baseline GSM8K and Human-eval results are cited from Wang et al. (2024b).

### v3

@@ Abstract
Efficient finetuning of large language models (LLMs) aims to adapt the LLMs with reduced computational and memory costs. Previous LoRA-based approaches initialize the low-rank matrices with Gaussian distribution and zero values while keeping the original weight matrices frozen. However, the trainable model parameters optimized in an unguided subspace might interfere with the well-learned subspace of the pretrained weight matrices. In this paper, we propose MiLoRA, a simple yet effective LLM finetuning approach that only updates the minor singular components of the weight matrix while keeping the principal singular components frozen. It is observed that the minor matrix corresponds to the noisy or long-tail information, while the principal matrix contains important knowledge. The MiLoRA initializes the low-rank matrices within a subspace that is orthogonal to the principal matrix, thus the pretrained knowledge is expected to be well preserved. During finetuning, MiLoRA makes the most use of the less-optimized subspace for learning the labeled dataset. Extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks present the superior performance of our method.
MiLoRA: Harnessing Minor Singular Components for
Parameter-Efficient LLM Finetuning
Hanqing Wang1,  Yixia Li2,  Shuo Wang3,  Guanhua Chen2thanks: Corresponding Authors. ,  Yun Chen1,411footnotemark: 1 1Shanghai University of Finance and Economics 2Southern University of Science and Technology,  3Tsinghua University 4MoE Key Laboratory of Interdisciplinary Research of Computation and Economics, Shanghai University of Finance and Economics
@@ 1 Introduction
Large language models (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023; Jiang et al., 2023, LLMs) have demonstrated superior performance on various tasks (Zheng et al., 2023), such as math reasoning (Wang et al., 2024b) and question answering (Ivison et al., 2023). These models are pretrained with the next token prediction task (Brown et al., 2020) on large web-scale data, then finetuned with instruction data as well as human preference data (Ouyang et al., 2022; Yu et al., 2023; Cui et al., 2023) for different downstream tasks. Fully finetuning is commonly employed to unlock the complete potential of LLMs, however, optimizing all model parameters necessitates substantial and restrictive computational resources (Touvron et al., 2023; Jiang et al., 2023), which hampers the utilization of LLMs across diverse scenarios.
@@
In this paper, we propose Minor singular component based Low Rank Adaptation (MiLoRA) for efficient LLM finetuning. MiLoRA has a similar framework as LoRA but employs a different initialization schedule. Specifically, a weight matrix ${\bm{W}}$ is decomposed with the singular value decomposition (SVD) algorithm. Based on the magnitude of the singular values, we divide ${\bm{W}}$ into two components: the principal matrix ${\bm{W}}_{p}$ corresponding to large singular values and the minor matrix ${\bm{W}}_{m}$ corresponding to small singular values. We argue that the principal matrix captures the essence of the pretrained knowledge, whereas the minor matrix is suboptimal with noisy or long-tail information. It is supported by previous works (Hajimolahoseini et al., 2021; Sharma et al., 2024; Li et al., 2024) that principal low-rank approximation can achieve comparable or even better performance to full parameters.
Motivated by these observations, we freeze the principal matrix ${\bm{W}}_{p}$ and adapt the minor singular components during finetuning. The low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ in LoRA framework are initialized with the minor matrix ${\bm{W}}_{m}$. Since the trainable low-rank matrices are initialized in a subspace orthogonal to the principal matrix, MiLoRA is expected to effectively learn from finetuning tasks while better preserving and utilizing the pretrained knowledge. To maintain the capability of the pretrained model at the start of finetuning, vanilla LoRA explicitly initializes ${\bm{B}}$ with zeros. In contrast, MiLoRA naturally satisfies this requirement, as the pretrained weight matrix ${\bm{W}}$ equals the frozen principal part ${\bm{W}}_{p}$ plus the low-rank part ${\bm{W}}_{m}={\bm{B}}{\bm{A}}$. We conduct extensive experiments on commonsense reasoning, math reasoning, instruction following and visual instruction following benchmarks. The experimental results show that MiLoRA consistently outperforms LoRA without sacrificing training or inference efficiency, such as commonsense reasoning (+1.6/+1.1 on LLaMA2-7B/LLaMA3-8B), math reasoning (+2.0 on LLaMA2-7B), instruction following (+2.9 on LLaMA2-7B) and visual instruction following (+1.4 on LLaVA1.5-7B). [^1] Our code and model are publicly available at https://github.com/sufenlp/MiLoRA.
@@ 2 Preliminaries
Given a matrix ${\bm{W}}\in\mathbb{R}^{m\times n}$, its singular value decomposition is denoted as ${\bm{W}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\top}$, where ${\bm{U}}=\left[u_{1},u_{2},\cdots,u_{m}\right]\in\mathbb{R}^{m\times m}$, ${\bm{V}}=\left[v_{1},v_{2},\cdots,v_{n}\right]\in\mathbb{R}^{n\times n}$. The columns of ${\bm{U}}$ are the left singular vectors, and the columns of ${\bm{V}}$ are the right singular vectors. The $\bm{\Sigma}\in\mathbb{R}^{m\times n}$ is a diagonal matrix containing the singular values of ${\bm{W}}$ in descending order. Without loss of generality, we suppose $m\leq n$ to simplify the notation. The SVD of ${\bm{W}}$ can be reformulated as
@@
where ${\bm{W}}^{(0)}$ and $\Delta{\bm{W}}$ refer to the pretrained weight and weight change, respectively. The $\alpha$ and $r$ are hyperparameters of scaling-factor and LoRA rank ($r\ll\min(m,n)$). During finetuning, the pretrained matrix ${\bm{W}}^{(0)}$ is kept frozen. It significantly diminishes the number of trainable parameters as both ${\bm{A}}$ and ${\bm{B}}$ matrices are low-rank. The ${\bm{B}}$ matrix is initialized with zero while ${\bm{A}}$ matrix adopts a random Gaussian distribution with zero mean value. This initialization strategy ensures the $\Delta{\bm{W}}=0$ at the beginning of training. The LoRA method only modifies the linear matrices in the Transformer model. The low-rank matrices can be easily merged into the pretrained linear matrix to get updated for inference, which does not require additional computing and GPU memory compared with full finetuning. However, the vanilla LoRA method fails to select the optimal subspace for updating the model parameters, as the low-rank matrices ${\bm{A}}$ and ${\bm{B}}$ are randomly initialized. This might potentially detract from the pretrained knowledge encoded in the pretrained weight matrix.
@@ 3 Methodology
The decomposition can also be reformulated in a matrix form. The ${\bm{U}}$ matrix in the SVD of ${\bm{W}}$ can be reformulated as ${\bm{U}}=\left[{\bm{U}}_{p},{\bm{U}}_{m}\right]$, where ${\bm{U}}_{p}=\left[u_{1},u_{2},\cdots,u_{m-r}\right]$ and ${\bm{U}}_{m}=\left[u_{m-r+1},u_{m-r+2},\cdots,u_{m}\right]$ are left singular vectors corresponding to principal and minor singular values, respectively. The ${\bm{V}}$ and ${\bm{\Sigma}}$ matrices can be reformulated similarly. The SVD of ${\bm{W}}$ can be expressed as:
@@
One potential variant of MiLoRA is to impose strict constraints on the orthogonality between ${\bm{W}}_{m}$ and ${\bm{W}}_{p}$ during fine-tuning. However, we adopt a softer approach by guiding the optimization direction solely through initialization, resulting in a method that is simpler, more training-efficient, and more flexible. Our experiment results show that our scheme works quite well in balancing learning from the finetuning dataset and preserving pretrained knowledge.
@@ 4 Experiments
Table 1: Commonsense reasoning evaluation results for LLaMA2-7B and LLaMA3-8B on eight tasks. The reported metric in this table is accuracy. Results are cited from Liu et al. (2024b). Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
@@
Table 2: Math reasoning evaluation results for LLaMA2-7B on GSM8K and MATH. The evaluation metric presented in this table is the Exact Match (EM) ratio. Results are cited from Yu et al. (2023). For each dataset, the highest performance scores among the different PEFT methods are shown in bold.
@@ 4.1 Experiments on Large Language Model
We compare MiLoRA with baselines on three different types of LLM downstream tasks.
[…]
Method AlpacaEval 2.0 FollowBench IFEval Avg.
LC WR WR Level1 Level2 Level3 Level4 Level5 Prompt-strict Ins-strict Prompt-loose Ins-loose
LoRA 5.6 3.9 46.5 39.2 35.5 22.2 14.3 28.8 39.9 31.1 42.3 28.1
PiSSA 5.6 3.9 44.2 37.1 30.6 20.9 14.3 31.6 42.7 34.6 45.7 28.3
MiLoRA 7.1 4.4 45.1 37.6 38.2 28.5 20.8 32.4 44.8 34.8 47.7 31.0
Table 3: Instruction following evaluation results for LLaMA2-7B on AlpacaEval 2.0, FollowBench, and IFEval. AlpacaEval 2.0 is evaluated using Length-controlled Win Rate (LC WR) and Win Rate (WR) against GPT-4 Turbo (OpenAI, 2023b), FollowBench adopts Hard Satisfaction Rate (HSR), and IFEval is measured by accuracy. Bold numbers indicate the highest performance scores for each dataset across the different PEFT methods for the corresponding model.
@@
We use the same hyperparameter configurations as Hu et al. (2023) without tuning for all methods. Details can be found in Appendix A.2. We denote this hyperparameter setup as LLM-Adapters. By default, we use a rank of 32. As math reasoning has a large training dataset, we set the rank to 64. We use the implementation of LLM-Adapters (Hu et al., 2023) [^2] https://github.com/AGI-Edgerunners/LLM-Adapters for commonsense reasoning, the implementation of PiSSA (Meng et al., 2024) [^3] https://github.com/GraphPKU/PiSSA for math reasoning and the implementation of open-instruct [^4] https://github.com/allenai/open-instruct for instruction tuning.
@@ Results
Table 4: Visual instruction tuning evaluation results for LLaVA1.5-7B on seven vision-language tasks. Accuracy serves as the metric reported in this table. We directly use checkpoints from Liu et al. (2024a) to reproduce the results of baseline methods.
We report results in Table 1 for commonsense reasoning, Table 2 for math reasoning, and Tables 3 for instruction following, respectively. For commonsense reasoning, we also include the ChatGPT baseline reported in Liu et al. (2024b) as a reference, which is obtained with GPT-3.5 Turbo (OpenAI, 2023a) API using a zero-shot Chain of Thought (Wei et al., 2022). For math reasoning, we add the full finetuning results from Yu et al. (2023) as a reference.
As can be seen, MiLoRA consistently surpasses all baseline methods across different datasets and LLMs, indicating that MiLoRA is a highly effective finetuning method. Specifically, for commonsense reasoning, MiLoRA outperforms LoRA and PiSSA by an average of 1.6 and 5.4 accuracy points on LLaMA2-7B and 1.1 and 6.5 points on LLaMA3-8B. For math reasoning, MiLoRA improves over LoRA and PiSSA by 2.0 and 3.7 averaged Exact Match (EM) points on LLaMA2-7B, respectively. However, it still underperforms full finetuning by an average of 2.5 EM score. This suggests that the PEFT methods still have room for improvement to fully match the performance of full fine-tuning. For instruction following, the results on AlpacaEval 2.0 indicate that PiSSA performs comparably to LoRA, while MiLoRA outperforms both by 1.5 in LC WR and 0.5 in WR, demonstrating its superior effectiveness. Furthermore, experiments on FollowBench and IFEval show that MiLoRA significantly surpasses both LoRA and PiSSA across these benchmarks. Notably, MiLoRA excels on more challenging problems (levels 4 and 5) in FollowBench, likely due to its ability to better retain pre-trained knowledge, which is crucial for handling complex instructions. Overall, MiLoRA surpasses LoRA and PiSSA in instruction following, achieving an average improvement of 2.9 and 2.7 points, respectively, further underscoring its advantages.
@@ 5.1 Comparison with More PEFTs
Table 5: Comparison with more baselines on math and code tasks. Baseline GSM8K and Human-eval results are cited from Wang et al. (2024c). The evaluation metric for GSM8K is Exact Match (EM), while for Human-eval, it is Pass@1.
To further demonstrate the effectiveness of MiLoRA, we compare MiLoRA with additional popular baselines, such as rsLoRA Kalajdzievski (2023), LoRA+ Hayou et al. (2024), DoRA Liu et al. (2024b), AdaLoRA Zhang et al. (2023), LoRA-GA Wang et al. (2024c). We follow the math and code experiment settings in Wang et al. (2024c). Specifically, we finetune LLaMA2-7B with MiLoRA on MetaMathQA 100K and Code-Feedback 100K for one epoch using rank $8$ and compare MiLoRA with baseline results reported in Wang et al. (2024c). Table 5 shows the results. We report the training cost of all methods relative to LoRA, which requires 3.58 hours of training on the math task using a single NVIDIA-L40 GPU. As can be seen, MiLoRA consistently outperforms all baselines, improving the best PEFT baseline LoRA-GA by 1.1 and 4.6 points on GSM8K and Human-eval, respectively. The training efficiency of MiLoRA is also very high. The time for SVD decomposition in MiLoRA is less than six minutes and negligible when compared with the total training time. Therefore, the training time required of MiLoRA is consistent with that of LoRA, while DoRA and AdaLoRA require additional training time overhead.